OpenStudy (anonymous):

find all values of theta 0

OpenStudy (anonymous):

Is this \[\cos2\theta~~or~~\cos ^2{\theta}\]?

OpenStudy (anonymous):



OpenStudy (anonymous):

Ok, that makes this problem much better. The double angle formula for cosine is \[\cos2\theta=1-2\sin ^{2}\theta\]So\[\sin \theta=1-2\sin ^{2}\theta\]set the equation to 0\[0=1-2\sin ^{2}\theta-\sin \theta\]Can you do it from here?

OpenStudy (anonymous):

no im still stuck on this wrstlr323